# SiCP: Simultaneous Individual and Cooperative Perception for 3D Object Detection in Connected and Automated Vehicles

### v1

@@ Abstract
Cooperative perception for connected and automated vehicles is traditionally achieved through the fusion of feature maps from two or more vehicles. However, the absence of feature maps shared from other vehicles can lead to a significant decline in object detection performance for cooperative perception models compared to standalone 3D detection models. This drawback impedes the adoption of cooperative perception as vehicle resources are often insufficient to concurrently employ two perception models. To tackle this issue, we present Simultaneous Individual and Cooperative Perception (SiCP), a generic framework that supports a wide range of the state-of-the-art standalone perception backbones and enhances them with a novel Dual-Perception Network (DP-Net) designed to facilitate both individual and cooperative perception. In addition to its lightweight nature with only 0.13M parameters, DP-Net is robust and retains crucial gradient information during feature map fusion. As demonstrated in a comprehensive evaluation on the OPV2V dataset, thanks to DP-Net, SiCP surpasses state-of-the-art cooperative perception solutions while preserving the performance of standalone perception solutions. The code link is https://github.com/DarrenQu/SiCP.
1 Introduction
Cooperative perception leverages wireless connectivity among vehicles to exchange sensor information, fostering a collective understanding of the driving environment. In contrast to individual perception, where a vehicle relies solely on local sensor data, cooperative perception creates a local map that extends beyond line-of-sight, enhancing vehicle navigation over an extended space and time horizon.
Within the realm of cooperative perception, feature map sharing-based approaches exhibit significant promise due to their superior performance and modest demands on the wireless channel capacity.
Figure 1: Comparison of LiDAR-based 3D object detection performance among different perception models. Except PointPillars [13], all others are cooperative perception models operating without over-the-air feature maps. The performance of individual perception mode is important as features shared by other vehicles may not be always available. Our SiCP model shows no performance degradation as compared to PointPillars [13].
Notable works in this domain, such as F-Cooper [2], AttFuse [26], Where2Comm [10] and V2X-ViT [25], have showcased the superiority of cooperative perception over its individual counterparts.
However, a critical issue overlooked by these prior works poses a challenge to the potential widespread adoption of their solutions. Specifically, they assume vehicles either exclusively use a cooperative perception model for both individual (in absence of feature maps from peers) and cooperative perception tasks or possess sufficient resources to concurrently employ two models, each dedicated to a different perception mode. Both assumptions present challenges, e.g., the former leads to performance degradation when the cooperative perception model operates in individual perception mode, as illustrated in Figure 1. Several cooperative perception models consistently lag behind PointPillars [13], the sole individual perception model, across all evaluated Intersection-Over-Union (IOU) thresholds in individual perception mode, with a detection precision gap reaching as high as 17.03%. Meanwhile, the latter assumption encounters practical hurdles, as vehicles often have limited hardware and software resources to support more than one perception solution. Existing cooperative perception models typically demand a larger memory footprint and computing power than individual perception solutions, posing obstacles to deployment. Hence, the community needs a cooperative solution that is lightweight and exhibits no performance degradation when operating in individual perception mode.
@@ 1 Introduction
Figure 2: Different approaches to perception. In (a), individual perception uses local sensor data for object detection. Cooperative perception, shown in (b), combines data from various vehicles to enhance the ego-vehicle’s perception. Simultaneous Individual and Cooperative Perception (SiCP), as depicted in (c), supports both functionalities simultaneously.
1.1 Proposed Solution
We introduce an innovative perception framework, SiCP (Simultaneous Individual and Cooperative Perception), offering a versatile and unified solution to concurrently support individual and cooperative perceptions. Illustrated in Figure 2 (c), SiCP comprises three key components: a feature extractor, a feature processor, and a detection head. This architecture is intentionally designed to align with popular pipelines of both individual perception models (Figure 2 (a)) and cooperative perception models (Figure 2 (b)). This alignment facilitates the seamless transformation of singular-function perception solutions into dual-mode perception systems.
More specifically, the feature extractor of SiCP generates features catering to both perception modes, storing them locally and transmitting them for cooperative perception at neighboring vehicles. To maintain manageable communication demands, we generate features based on the sender vehicle’s perspective, eliminating the need for the sender vehicle to compute and share features individually with each nearby receiver as in  [15]. An initial affine transformation in the feature processor aligns the received features’ perspective with the receiver vehicle’s.
SiCP’s feature processor orchestrates local and received features to optimize object information representation. Housing a novel Dual-Perception network (DP-Net), it acts as a conduit for local features in individual perception mode while efficiently performing feature fusion in the presence of received features. This fusion component effectively preserves and combines gradient information from both local and received features, enabling the detection of objects not recognizable using local or received features alone in cooperative perception mode. Further details about the feature processor are provided in Section 3.1.
In SiCP’s detection head, we adopt a standard architecture as seen in [13]. However, we tailor its loss function to accommodate losses from both individual and cooperative perceptions. Special attention is given to the training strategy to balance the loss ratio from each perception mode, facilitating the feature extractor in capturing patterns useful for both modes. Our performance evaluation shows that SiCP excels in both individual and cooperative perception.
It’s important to note that unlike other cooperative perception models  [26, 25], SiCP fuses feature maps from only two vehicles in a single forward propagation. This design choice is motivated by the reality that a vehicle has to wait longer to accumulate feature maps from more neighbors. Given the lossy nature of wireless communication, it is uncertain if a feature map from a third surrounding vehicle will ever arrive. To counteract delays that offset the benefits of cooperative perception, SiCP focuses on fusing feature maps from one vehicle at a time. Different neighboring vehicles’ feature maps can be selected at different times, especially when multiple connected and automated vehicles are available.
1.2 Main Contributions
@@
  • For the first time, we recognize the significance of simultaneous individual and cooperative perception in connected and automated vehicles.
  • The proposed solution excels in cooperative perception task, surpassing the state-of-the-art models in performance, while achieving result on par with leading standalone 3D detection models in individual perception task.
  • The proposed DP-Net is a Plug-and-Play module as it can be seamlessly integrated into other individual perception models, enabling simultaneous individual and cooperative perception.
  • The proposed DP-Net is also a lightweight component, comprising just 0.13M parameters, representing a mere 1.7% increase from the PointPillars [13] backbone model.
  • The proposed SiCP framework will facilitate the exploration of novel and effective designs for feature extractors, feature fusion, and detection heads, enabling simultaneous individual and cooperative perception.
2 Related Works
2.1 Individual Perception
@@
2.2 Cooperative Perception
Cooperative perception solutions for connected and automated vehicles can be classified into early fusion [3, 1], deep fusion [2, 26, 22, 25, 4, 14, 15, 9, 19, 21, 29], and late fusion [30, 8, 6, 5]. Among these, the deep fusion method strikes a balance between bandwidth and detection performance, making it widely embraced in the literature.
Activation function based fusion. Initially introduced in [2], F-Cooper employs the maxout operation for feature map fusion. CoFF [9] enhances F-Cooper by incorporating feature enhancement techniques. Despite their advancements, these solutions grapple with challenges stemming from the heterogeneity of feature maps originating from different vehicles. Even when focusing on the same region, perceptual differences can result in significantly varied features.
Attention based fusion. AttFuse [26] incorporates a self-attention operation for feature fusion. Despite its effectiveness, the solution overlooks nearby features, missing out on crucial information locality. V2X-ViT [25] introduces a unified transformer architecture for heterogeneous multi-agent perception, while COOPERNAUT [4] adopts a transformer-based architecture with point-based representations. While these solutions consider self-attention relations among all points in feature maps, making them computationally intensive, they are less focused on specific regions.
Other models adopt diverse strategies for feature fusion from various perspectives. For instance, Where2comm [10] introduces a spatial confidence map to capture the spatial diversity of perceptual data, effectively minimizing communication bandwidth. V2VNet [22] employs a variational image compression algorithm to compress features and utilizes the mean operation for feature aggregation. DiscoNet [14] employs a teacher-student distillation method that combines early and deep fusion techniques. Additionally, CoAlign [15] introduces a novel hybrid collaboration framework designed to address pose errors. Furthermore, Camera-LiDAR frameworks enhance cooperative perception with camera support [23].
@@ 2.2 Cooperative Perception
Figure 3: An overview of the SiCP architecture showcases its components: a feature extractor, a feature processor (DP-Net), and a detection head. All vehicles have identical feature extractors producing fusible features. The feature processor manages local features $F_{ego}$ for individual perception and fused features $F^{*}_{ego}$ for cooperative perception. Features from other vehicles (e.g., $F_{j}$) are transformed to the ego vehicle’s perspective (as shown in (a)) and integrated with local features using the DP-Net module (as shown in (b)). The resulting feature $F^{*}_{ego}$ is then processed by the detection head to generate classification and regression results for either individual or cooperative perceptions.
3 Methodology
@@
This section outlines the major components of the SiCP (Simultaneous Individual and Cooperative Perception) framework, illustrated in Figure 3. The framework comprises three key elements: a feature extractor, an Dual-Perception Network, and a unified detection head. We adopt the backbone of PointPillars [13] as our feature extractor, aiming to achieve an optimal balance between effectiveness and efficiency in 3D object detection. We propose the Dual-Perception Network (DP-Net) that seamlessly integrates with the existing backbone (Section 3.1), adeptly processing local features for individual perception and fusing features from multiple vehicles for cooperative perception. We employ a unified detection head (Section 3.2) to process features in both individual and cooperative perception tasks. The proposed solution operates on the premise of vehicular trustworthiness, with all vehicles employing the same machine-learning model for executing their object detection tasks.
3.1 Dual-Perception Network (DP-Net)
@@
The proposed DP-Net module should ensure that the individual perception process and cooperative perception process run in parallel. Moreover, the resulting feature maps, whether for individual or cooperative perception, should be compatible with each other and capable of being processed by a unified detection head. Specifically, DP-Net executes operations based on whether the ego vehicle has received features from neighboring vehicles. In the absence of received features, DP-Net continues utilizing the current ego vehicle’s feature maps, thereby guaranteeing its individual perception performance. Upon receiving features from neighboring vehicles, DP-Net performs a perspective transformation on these features before forwarding them to the fusion module for integration. We will delve into the critical endeavor of effectively fusing feature maps in Sections 3.1.1 and 3.1.2.
3.1.1 Receiver-Agnostic Feature Sharing
@@
To enable cooperative perception through deep fusion, vehicles must share their locally generated features with nearby vehicles. Achieving efficient vehicular communications requires implementing a receiver-agnostic feature-sharing approach, wherein the sender does not need to know the location and pose of the potential receivers. However, implementation of existing methods often require transforming LiDAR data or feature maps into the perspective of the receiving vehicle before transmission [26, 25, 29, 21, 19]. In scenarios involving multiple receiving vehicles, this results in creating and transmitting multiple versions of the same feature maps, leading to significant network traffic and computational demands on the sending vehicle. Alternatively, the sending vehicle can optimize its communication by broadcasting its local feature map to all nearby vehicles. In this approach, the sending vehicle shares not just its feature map but also its current location and pose information. Upon receiving a shared feature map, the recipient vehicle performs feature transformation using the Affine Transformation technique [15].
3.1.2 Complementary Feature Fusion
@@
Differences in the LiDAR point cloud data acquired by different vehicles can result in varying features for the same object/vehicle. Due to occlusions, the receiver has difficulty in capturing meaningful features for one object, indicated by the red box in Figure 4 (b). Consequently, the resulting feature maps might misinterpret these regions as background, resulting in relatively larger numerical representations. When the feature maps from the sender and receiver are fused using the basic maxout function, the gradients in the sender’s features vanish due to the influence of the larger numbers in the receiver’s feature map.
@@
Next, we normalize $\tilde{M}$ to get a normalized weighted map $M$, in which all the numbers range within $[0,1]$. As $M$ is used to weigh and fuse features from two feature maps, we further adjust $M$ as follows. For any element $m_{ij}$ in $M$, we have
@@
As such, the weight map $M$ only provides clues to fuse the features within the overlapping area between $\mathcal{F}_{ego}$ and $\mathcal{F}_{j}^{{}^{\prime}}$. While for the non-overlapping area, the weight map is always 00. This implies the ego vehicle does not consider other vehicles’ data but only relies on its own to detect objects.
@@
where $\otimes$ denotes the element-wise multiplication. Note that the weight map $M$ is responsible for adjusting the ego vehicle’s location feature map, while the complementary weight (CW) map $(1-M)$ modifies the received feature map. This means, in the fused feature map, each point is strongly influenced either by the ego vehicle’s feature or by the other vehicle’s feature, but not both simultaneously.
3.2 Detection Head
@@
To realized effective training, we need to ensure an equitable distribution of training data for individual and cooperative perception tasks. Due to resource limitations, conducting additional training rounds is unfeasible. Both individual and cooperative perception components in the proposed model must be trained using a single input data and labels. Therefore, the total number of training rounds remains constant. Specifically, our methodology involves inputting a single data instance, encompassing the raw local LiDAR data and a feature map shared by another vehicle, into the network. Then, backpropagations are applied, allowing the individual perception and cooperative perception pipelines to be trained jointly.
4 Experiments
Datasets. We conduct our evaluations using the OPV2V dataset [26] which is a simulation dataset for V2V cooperative perception, co-simulated through OpenCDA [24] and CARLA [7]. In accordance with the particular training and testing specifications of our model, we implement a First-Come-First-Serve policy for each frame in the dataset. This approach facilitates the utilization of data from two vehicles: the ego vehicle and the sender vehicle. Following [26], the LiDAR detection range of each vehicle to be $x\in[-140.8m,140.8m]$ and $y\in[-40m,40m]$, and the communication range between vehicles is $70m$.

### v2

@@ Abstract
Cooperative perception for connected and automated vehicles is traditionally achieved through the fusion of feature maps from two or more vehicles. However, the absence of feature maps shared from other vehicles can lead to a significant decline in 3D object detection performance for cooperative perception models compared to standalone 3D detection models. This drawback impedes the adoption of cooperative perception as vehicle resources are often insufficient to concurrently employ two perception models. To tackle this issue, we present Simultaneous Individual and Cooperative Perception (SiCP), a generic framework that supports a wide range of the state-of-the-art standalone perception backbones and enhances them with a novel Dual-Perception Network (DP-Net) designed to facilitate both individual and cooperative perception. In addition to its lightweight nature with only 0.13M parameters, DP-Net is robust and retains crucial gradient information during feature map fusion. As demonstrated in a comprehensive evaluation on the V2V4Real and OPV2V datasets, thanks to DP-Net, SiCP surpasses state-of-the-art cooperative perception solutions while preserving the performance of standalone perception solutions. The source code can be found at https://github.com/DarrenQu/SiCP.
I Introduction
Automated vehicles rely on standalone perception models to detect and comprehend 3D objects in their surrounding environment through individual sensors. Cooperative perception, on the other hand, allows multiple vehicles to collaborate, enhancing their collective environmental awareness. Traditionally, these two modes of perception were studied in isolation, as shown in Figure 1 (a) and (b), neglecting the synergies between them. The significance of concurrently addressing these two facets was undervalued.
In the realm of cooperative perception, researchers commonly adapt solutions from individual perception to collaborative settings. This adaptation involves expanding popular standalone 3D detection models, such as PointPillars [1], SECOND [2] and VoxelNet [3], by fusing feature maps to become a cooperative perception models like F-Cooper [4], AttFuse [5], V2X-ViT [6] and CoBEVT [7]. With the exception of F-Cooper, most alternative methods require the sender to modify its local feature data to match the receiver’s viewpoint prior to transmission, which proves to be impractical. The impracticality arises due to the potential existence of multiple receivers, and executing numerous transformations and transmissions of local feature maps introduces substantial computational and network overhead,
@@ I Introduction
Figure 1: Different approaches to 3D perception. In (a), individual perception uses local sensor data for object detection. Cooperative perception, shown in (b), combines data from various vehicles to enhance the ego-vehicle’s perception. Simultaneous Individual and Cooperative Perception (SiCP), as depicted in (c), supports both functionalities simultaneously.
making the entire process excessively costly. More importantly, the core problem with these cooperative perception models is their inability to effectively handle individual perception tasks. This concern is critical because, in real-world scenarios, there may not always be a sender vehicle available to share perception information. In such cases, ego vehicle must rely solely on its own perception capabilities to understand its environment. Our experiments indicate that cooperative perception models, as presented in Figure 1 (b), significantly lag behind the standalone perception model, as shown in Figure 1 (a), in performance on individual perception tasks.
This performance gap highlights the limitations of cooperative perception models in object detection scenarios where data is not shared between vehicles. A fundamental challenge arises from the divergence in features extracted from data of individual vehicles and multiple cooperating vehicles. The impact of these disparities remains poorly understood. To bridge this gap and pave the way for more robust and accurate automated vehicle perception systems, we explore the possibility of combining individual and cooperative perception to develop a unified framework that seamlessly integrates both individual and cooperative perceptions for connected and automated vehicles.
I-A Proposed Solution
We present a pioneering perception framework, called SiCP (Simultaneous Individual and Cooperative Perception), to handle both individual and cooperative perceptions simultaneously, as illustrated in Figure 1 (c). The SiCP architecture is composed of three key components: a feature extractor, a feature processor, and a detection head. Initially, each vehicle utilizes an identical feature extractor to generate features suitable for both individual perception and cooperative perception tasks. We devise a novel feature processor, named the Dual-Perception Network (DP-Net), to proficiently manage local features for individual perception and integrate fused features from neighboring vehicles for cooperative perception. In situations where features from neighboring vehicles are unavailable, DP-Net relies on the ego vehicle’s local feature map for individual perception tasks. When such features are accessible, they are first transformed to the ego vehicle’s perspective and then fused with ego vehicle’s local feature map. The DP-Net effectively merges Bird’s Eye View (BEV) feature maps by concatenating and condensing them into a single-channel feature map. This condensed map undergoes further processing through two convolutional layers, resulting in a weighted map. This weighted map plays a crucial role in adjusting the ego vehicle’s local feature map, while its complementary counterpart modifies the feature map received from other vehicles. The adjusted feature maps are then concatenated and reshaped to the desired output size, successfully integrating information from neighboring vehicles in cooperative perception scenarios.
In terms of practical implementation, we embrace an approach akin to First-Come-First-Serve (FCFS), wherein the feature map of the ego vehicle is fused with those of the initially received neighbors, eliminating the necessity to await additional features before initiating the fusion process. The resulting fused feature set is fed into the detection head, where it undergoes processing to generate classification and regression results for cooperative perception tasks. In cases where no additional features are received, the original local feature map is processed by the same detection head to complete individual perception tasks. This approach ensures seamless integration of individual and cooperative perception, thereby improving overall efficiency and accuracy of the connected and automated vehicle perception system.
I-B Main Contributions
@@
  • For the first time, we recognize the significance of employing a single 3D object detection network for simultaneous individual and cooperative perception within connected and automated vehicles. We present the SiCP framework as a novel solution to fill this gap.
  • The proposed DP-Net is an innovative Plug-and-Play module as it can be seamlessly integrated into other standalone 3D detection models, enabling simultaneous individual and cooperative perception.
  • The proposed DP-Net is also a lightweight component, comprising just 0.13M parameters, representing a mere 1.7% increase from the standalone 3D perception model [1].
  • The proposed DP-Net exhibits robustness in addressing alignment errors caused by asynchronous communication and inaccurate localization between vehicles.
II Related Works
II-A Individual Perception
@@
II-B Cooperative Perception
Cooperative perception solutions for connected and automated vehicles classified into early fusion [11, 12], deep fusion [4, 5, 6, 7, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23], and late fusion. Among these, the deep fusion strikes a balance between bandwidth and detection performance, making it widely embraced in the literature.
Activation function based deep fusion. Initially introduced in [4], F-Cooper employs the maxout operation for feature map fusion. CoFF [15] enhances F-Cooper by incorporating feature enhancement techniques. Despite advancements, these solutions grapple with challenges stemming from heterogeneity of feature maps originating from different vehicles. Even when focusing on the same region, perceptual differences can result in significantly varied features.
Attention based deep fusion. AttFuse [5] incorporates a self-attention operation for feature fusion. Despite its effectiveness, the solution overlooks nearby features, missing out on crucial information locality. V2X-ViT [6] introduces a unified transformer architecture for heterogeneous multi-agent perception, while CoBEVT [7] presents a generic transformer-based framework.
Figure 2: An overview of the SiCP architecture showcases its components: a feature extractor, a feature processor (DP-Net), and a detection head. All vehicles have identical feature extractors producing fusible features. The feature processor manages local features $F_{ego}$ for individual perception and fused features $F^{*}_{ego}$ for cooperative perception. Features from other vehicles (e.g., $F_{j}$) are transformed to the ego vehicle’s perspective and then performs a complementary fusion with the local features of the ego vehicle. The resulting feature $F^{*}_{ego}$ is then processed by the detection head to generate classification and regression results for either individual or cooperative perceptions.
While these solutions consider self-attention relations among all points in feature maps, making them computationally intensive, they are less focused on specific regions. Other models adopt diverse strategies for feature fusion from various perspectives. For instance, Where2comm [13] introduces a spatial confidence map to capture the spatial diversity of perceptual data, effectively minimizing communication bandwidth. Additionally, CoAlign [14] introduces a novel hybrid collaboration framework designed to address pose errors. Furthermore, Camera-LiDAR frameworks enhance cooperative perception with camera support [18].
@@ II-B Cooperative Perception
III Methodology
@@
This section outlines the major components of the SiCP (Simultaneous Individual and Cooperative Perception) framework, illustrated in Figure 2. The framework comprises three key elements: a feature extractor, a Dual-Perception Network, and an unified detection head. We adopt the backbone of PointPillars [1] as our feature extractor, aiming for an optimal balance between effectiveness and efficiency in 3D object detection. We propose the Dual-Perception Network (DP-Net) that seamlessly integrates with the existing backbone (Section III-A), adeptly processing local features for individual perception and fusing features from multiple vehicles for cooperative perception. We employ a unified detection head (Section III-B) to process features for both individual and cooperative perception tasks. The proposed solution operates on the premise of vehicular trustworthiness, with all vehicles employing the same machine-learning model for executing their object detection tasks.
III-A Dual-Perception Network (DP-Net)
@@
The proposed DP-Net module ensures that the individual perception process and cooperative perception process run in parallel. Moreover, the resulting feature maps, whether for individual or cooperative perception, should be compatible with each other and capable of being processed by a unified detection head. Specifically, DP-Net executes operations based on whether the ego vehicle has received features from neighboring vehicles. In the absence of received features, DP-Net continues utilizing current ego vehicle’s feature maps, thereby guaranteeing its individual perception performance. Upon receiving features from neighboring vehicles, DP-Net performs a perspective transformation on these features before forwarding them to the fusion module for integration. We will delve into the critical endeavor of effectively fusing feature maps in Sections III-A1 and III-A2.
III-A1 Receiver-Agnostic Feature Sharing
@@
To enable cooperative perception through deep fusion, vehicles must share their locally generated features with nearby vehicles. Achieving efficient vehicular communications requires implementing a receiver-agnostic feature-sharing approach, wherein the sender does not need to know the location and pose of the potential receivers. However, implementation of existing methods often require transforming LiDAR data or feature maps into the perspective of the receiving vehicle before transmission [5, 6, 24, 25, 26]. In scenarios involving multiple receiving vehicles, this results in creating and transmitting multiple versions of the same feature maps, leading to significant network traffic and computational demands on the sending vehicle. Alternatively, the sending vehicle can optimize its communication by broadcasting its local feature map to all nearby vehicles.
@@
In this approach, the sending vehicle shares not just its feature map but also its current location and pose information. Upon receiving a shared feature map, the recipient vehicle performs feature transformation using the Affine Transformation technique [14].
III-A2 Complementary Feature Fusion
@@
Differences in the LiDAR point cloud data acquired by different vehicles can result in varying features for the same object/vehicle. Due to occlusions, the receiver has difficulty in capturing meaningful features for one object, indicated by the red box in Figure 4 (b). Consequently, the resulting feature maps might misinterpret these regions as background, resulting in relatively larger numerical representations. When the feature maps from the sender and receiver are fused using F-Cooper [4] method (maxout function), the gradients in the sender’s features vanish due to the influence of the larger numbers in the receiver’s feature map.
@@
Next, we normalize $\tilde{M}$ to get a normalized weighted map $M$, in which all the numbers range within $[0,1]$. As $M$ is used to weigh and fuse features from two feature maps, we adjust $M$ as follows. For any element $m_{ij}$ in $M$, we have
@@
As such, the weight map $M$ only provides clues to fuse features within the overlapping area between $\mathcal{F}_{ego}$ and $\mathcal{F}_{j}^{{}^{\prime}}$. While for the non-overlapping area, the weight map is always 00. This implies the ego vehicle does not consider other vehicles’ data but only relies on its own to detect objects.
@@
where $\otimes$ denotes the element-wise multiplication. Note that the weight map $M$ is responsible for adjusting the ego vehicle’s location feature map, while the complementary weight map $(1-M)$ modifies the received feature map. This means, in the fused feature map, each point is strongly influenced either by the ego vehicle’s feature or by the other vehicle’s feature, but not both simultaneously.
III-B Detection Head
@@
To realized effective training, we need to ensure an equitable distribution of training data for individual and cooperative perception tasks. To conserve computational resources, conducting additional training rounds is unfeasible. Both individual and cooperative perception components in the proposed model must be trained using a single input data and labels. Therefore, the total number of training rounds remains constant. Specifically, our methodology involves inputting a single data instance, encompassing the raw local LiDAR data and a feature map shared by another vehicle, into the network. Then, backpropagations are applied, allowing the individual perception and cooperative perception pipelines to be trained jointly.
IV Experiments
Datasets. We conduct our evaluations on two extensively used datasets: V2V4Real [29] and OPV2V [5]. Both supporting Vehicle-to-Vehicle (V2V) cooperative perception research by providing numerous annotated scenes to facilitate algorithm development and evaluation. V2V4Real is a large-scale real-world dataset, collected by two vehicles simultaneously in the same location, providing multi-view sensor datastream. OPV2V is a simulation dataset featuring a variety of virtual cities and environments, co-simulated through OpenCDA [30] and CARLA [31].
In accordance with the particular training and testing specifications of our model, we implement a First-Come-First-Serve policy for each frame in both datasets. This approach facilitates the utilization of data from two vehicles: the ego vehicle and the sender vehicle.